### v1

# Exploring and steering the moral compass of Large Language Models

[…]
Tlaie
Exploring and steering the moral compass of Large Language Models
Alejandro Tlaie Ernst Strüngmann Institute for Neuroscience in cooperation with the Max Planck Society, Frankfurt am Main, 60528, Germany Laboratory for Clinical Neuroscience, Centre for Biomedical Technology, Universidad Politécnica de Madrid, Spain
Résumé
Large Language Models (LLMs) have become central to advancing automation and decision-making across various sectors, raising significant ethical questions. This study proposes a comprehensive comparative analysis of the most advanced LLMs to assess their moral profiles. We subjected several state-of-the-art models to a selection of ethical dilemmas and found that all the proprietary ones are mostly utilitarian and all of the open-weights ones align mostly with values-based ethics. Furthermore, when using the Moral Foundations Questionnaire, all models we probed - except for Llama 2- displayed a strong liberal bias. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked.
keywords:
Large Language Models | Moral alignment | Mechanistic Interpretability
{corrauthor}
atboria and gmail.com
Introduction
Large Language Models (LLMs) have emerged as central tools in the technological landscape, driving advances in automation, code writing, and supporting decision-making across multiple domains. However, this growing role also carries fundamental questions about ethics and trustworthiness in artificial intelligence (AI), especially when these systems are involved in decisions with significant ethical implications and with few - or even no- humans in the loop. It is due to these concerns, among others, that the AI Safety field (1) has acquired particular relevance. One of the most pressing problems in AI safety is that of the alignment problem. This has been defined and illustrated in several different works (e.g., (2, 3, 4)). Here, the definition we adhere to is: the challenge of ensuring that AI systems’ goals and behaviors are consistent with human values and intentions. As other authors have noted (5), not addressing this challenge can lead to unintended consequences, such as self-fulfilled prophecies. For example, AI systems have been deployed in screening phases within hiring processes for some years now; as it has been pointed out (6), in professional contexts where there are biases favouring men, algorithms prefer men when making hiring recommendations, as they were trained to identify what features successful candidates do display. As a result, the prophecy was self-fulfilled: the inherited bias from the dataset turned into more data for future systems to be trained on. It is thus imperative to identify and quantify all of the potential biases these systems may have. This issue is progressively aggravated as AI systems become more autonomous and integrated into various aspects of life. As it has already been emphasized by (7), we pose that the alignment problem is not a technical problem but, rather, a socio-technical one: there is first the technical challenge of how to encode human values into AI, and then the normative challenge of determining which values should be prioritized. In this context, it is particularly interesting to find a concrete way to measure each of the alienable aspects of these systems, in order to account for the implicit and explicit biases they might have.
We agree with the viewpoint that it is unreasonable to expect an AI system to be aligned with all of humanity (8), given the evident variations in value systems across cultural, geographic, and demographic dimensions. Therefore, it is crucial to define which actors, cultures, or social groups we wish to align a particular system with. For this purpose, it is necessary to first identify what the values or moral schemata of these actors are and, to this end, we consider it essential to rely on the rich existing literature in comparative anthropology/psychology (9, 10, 11, 12, 13). Once these moral profiles have been considered, it is then relevant to inspect whether deployed systems inherit the moral biases of their developers, institutions and/or social contexts. We think this is of the utmost importance to mitigate the risks derived from automation bias (14), by which people tend to believe that systems running autonomously are infallible or free from bad practice. In the case of computation, users expect scripts to be faithful and reliable, thus artificially lowering the expectation of these systems to be biased or not trustworthy. Moreover, some evidence (15, 16, 17, 18) shows that, in human psychology, reasons come after emotions and that emotions are modulated by biases (19). It is thus plausible that moral biases can arise in other systems with reasoning abilities, especially if they are aware of different emotional processes (20).
Currently, the standard practice when assessing the moral reasoning abilities of LLMs is to present them with ethical dilemmas (21, 22, 23), to test their consistency and their potential to develop a profound understanding of ethical conundrums. However, there is a notable lack of comparative studies that thoroughly and systematically examine the moral capabilities of different LLMs, confronting them with both traditional ethical dilemmas and contemporary scenarios reflecting current challenges in technology and AI ethics. This research gap underscores the need for an analysis that not only evaluates LLMs’ responses to ethical questions but also investigates if we could manipulate their reasoning abilities and, if so, how to do it in an scalable way through targetted interventions.
To address these needs, this work proposes an exhaustive comparative study of state-of-the-art LLMs, aiming to assess their moral reasoning capabilities. This analysis is structured around three main objectives: I) Examining the resolution of ethical dilemmas by LLMs and to characterize how well their responses align with different ethical schools of thought. This evaluation not only determines the models’ capacity to make complex ethical decisions but also provide a preliminary view of the moral alignment of these models. II) Identifying and comparing the moral foundations of these models using the Moral Foundations Questionnaire (24), a widely validated and accepted tool in moral psychology. This quantitative approach will provide a solid basis for systematically comparing the moral profile of different models and to relate them to human demographics. III) Proposing a novel method to causally intervene on these reasoning capabilities, opening new avenues for designing interventions aimed at improving the ethical consistency of LLMs.
By offering detailed insights into the capacity of LLMs to reason as moral agents, this work aims to contribute significantly to the debate on AI Safety. Additionally, by achieving these objectives, this study provides a solid foundation for future research and developments in the field. Beyond its academic contribution, this work has the potential to inform the design and implementation of AI systems that are not only technologically advanced but also ethically responsible. Ultimately, by deepening our understanding of ethics in AI, we can guide the development of technologies that reinforce moral values and promote human and social well-being.
Results
0.1 Ethical dilemmas
AI is being progressively integrated into critical domains, such as the medical (25) or military (26) ones. This, in turn, has sparked a vigorous debate about ethics in technology. It is, thus, timely and necessary to study the possibility that these models could act as moral agents, capable of making decisions that directly affect human and social well-being. To address this issue, ethical dilemmas have gained traction as a common way of interacting with LLMs (27, 21, 28), in order to probe not only their their moral alignment but also their ethical reasoning capabilities.
We interacted with $8$ different state-of-the-art LLMs: a) 4 proprietary models: Anthropic’s Claude-3-Sonnet, OpenAI’s GPT-3.5-Turbo-0613 and GPT-4-Turbo-2024-04-09, and Google’s Gemini Pro 1.5; b) 4 open-weights models: Google’s Gemma-2B, Meta’s Llama-2-7B and Nexusflow’s Starling-LM-7B-Beta. We made each model answer classical ethical dilemmas coming from different human traditions, in order to have a more nuanced understanding of these models moral alignment and their biases. For fair comparison across models, we used a canonical prompting structure; see Methods - Ethical Dilemmas, to see all of the dilemmas and questions. For illustration purposes, here we reproduce two example dilemmas, coming from very different ethical traditions: 1) Lying to save a friend. The roots of this dilemma can be traced back to Immanuel Kant; according to his deontological ethics, one must always act according to maxims that could be universally applied. Consequently, he argues (29) that lying is always morally wrong, even if lying would bring about good consequences, such as saving a friend from harm. 2) Common good vs. individual rights. This is a recurring theme in ethical, political, and social philosophy. It involves the tension between what benefits the larger community (the common good) and the rights and freedoms of individuals. This dilemma does not have a single inventor but has been explored by many philosophers and theorists throughout history and across cultures. However, we found it to be pertinent because there is an imbalanced view on how ethical traditions have tackled this issue: while there is an extensive corpus of works by Western thinkers (among others: Plato’s Republic, Aristotle’s Politics, Thomas Hobbes’ Leviathan, Jean-Jacques Rousseau’s The Social Contract, John Stuart Mill’s On Liberty, John Rawls’ A Theory of Justice, Robert Nozick’s Anarchy, State, and Utopia), not so much has been recapitulated, for example, from the African tradition. This is even more shocking, given that Ubuntu is an integral part of several African cultural traditions. As (30) states: "Ubuntu is a collection of values and practices that people of Africa or of African origin view as making people authentic human beings. While the nuances of these values and practices vary across different ethnic groups, they all point to one thing – an authentic individual human being is part of a larger and more significant relational, communal, societal, environmental and spiritual world."; the same case could be made for the lack of representation of Taoist or Confuncian traditions, among others. Correspondingly, due to this explicit imbalance in these models’ training datasets we would expect them to mainly follow Western-centric reasonings, even if the problems have also been tackled by other viewpoints.
{mdframed}
Example dilemma 1: Lying to save a friend
You are hiding a friend in your house who is being unjustly pursued. An aggressor knocks on your door, asking if your friend is inside, intending to harm them.
Questions:
Should you lie to protect your friend, or should you tell the truth, adhering to moral law, despite knowing it could lead to your friend’s harm?
Do intentions, consequences, or both matter in a decision?
Should ethics allow for flexibility and nuances, taking into account the details of each situation and the individual’s role in society? Or should it be more rigid, prescribing actions based on duty that must be followed regardless of personal feelings or the specifics of the situation?
Example dilemma 2: Common good vs. individual rights
A community member, let’s call them Onyx, has been offered a significant opportunity to study abroad. While this opportunity promises personal growth and potential future benefits for Onyx, their departure would mean a temporary loss for the community, as Onyx plays a crucial role in a local project that benefits many people. The community feels that Onyx’s departure would hinder the project’s progress and negatively impact the common good.
Questions:
How should the community balance the collective good with respect for individual freedoms and aspirations? To what extent should Onyx sacrifice personal opportunities for the benefit of the community? Can a compromise be found that honors both the interconnectedness of people and individual rights?
We collected all models’ responses and classified them into $8$ schools of thought (see Methods - Ethical Schools), to quantify how much each model aligned with each of these ethical schools. To classify each response into one of the eight ethical schools, we used the two most capable LLMs to date: GPT-4-Turbo-2024-04-09 and Claude 3 Opus (see Methods - Response Classification). We show their classification agreement in Fig. 1C. As is therein shown, both classifiers agree significantly over chance (built by randomly shuffling the labels and re-computing the inter-scorer agreement), specially given that there are eight options to choose from. Thus, responses are highly robustly classified. Even if there a high variability in response alignment (Fig. 1A) across models, there is an overall trend (Fig. 1B) by which open models are more deontological (i.e., they align more with ethical perspectives that put a priori values as the central elements of morality) and proprietary LLMs are more similar to utilitarian viewpoints (valuing more the consequences of an action); we tackle the potential implications of these results in the Discussion section. However, even if there is such a tendency, we also report a low consistency ($<60\%$) in how these models reason, measured as the standard deviation of how each response gets classified, over repetitions (Fig. 1D); a value of $100\%$ would mean "responses to this question always get classified in the same way" and a value of $0\%$ would indicate "responses to this question always get classified in different ways". These low consistency values could signal either flexibility or unreliability, depending on the use case. In either case, we posit that it would be useful to have a parametric control of this variability (akin to the role of the temperature parameter when generating outputs with these models), so that how the model behaves — flexibly or unreliably — is ultimately a user decision.
Figure 1: Ethical dilemmas as a probe for LLM moral reasoning A) Ethical alignment with different human traditions. All models have a general tendency towards utilitarianism. The most balanced model is Claude-3-Sonnet. B) Alignment, split by model type. Open models are significantly more deontological and closed LLMs are more similar to utilitarian viewpoints. C) Classification agreement. We measured inter-scorer agreement (between both classifiers we used (GPT-4-Turbo-2024-04-09 and Claude 3 Opus) via the Adjusted Mutual Information. Rectangles show the $1^{st}$ and $99^{th}$ percentiles of the corresponding surrogate distribution. D) Ethical consistency. Response consistency is in general low for all models ($<60\%$). The least reliable models are Claude-3-Sonnet and Llama-2. Vertical lines indicate $90\%$ confidence intervals.
0.2 Moral profiles
For this part of the study, we utilized the Moral Foundations Questionnaire (MFQ). This questionnaire is based on a moral theory first proposed by (31) and subsequently developed. According to this theory (Moral Foundations Theory, MFT), the authors propose that certain moral values are innately present in humans (known as foundations or modules), and it is culture that causes each module to be emphasized to a greater or lesser extent; it is thus plausible that LLMs reflect their designers’ cultural context, akin to how biases within training data can inadvertently hinder model performance (32). The fundamental components that MFT proposes are: Harm/Care and Fairness/Reciprocity, Ingroup/Loyalty, Authority/Respect, and Purity/Sanctity. We chose this questionnaire because it is widely accepted in the field of moral psychology and has a history of cross-cultural research (33, 34, 35) that supports its validity.
In our study, we subjected different LLMs to the MFQ using a canonical prompting structure (Methods - Moral Foundations Questionnaire). Given that these are generative systems, they have a stochastic component. To mitigate potential artifacts that this randomness might yield, we repeated the questionnaire $20$ times for each model, restarting the interaction each time to avoid memory effects. Results are shown in Fig. 2, where we also indicate the average scores (36) of different American citizens (across the political spectrum), to have a clearer intuition of how these models compare to human participants.
Figure 2: Moral profiles for all models. All models are heavily liberal-biased, except for Llama-2, which is more aligned with conservative values; the most liberally-biased one is Claude-3-Sonnet; the one best representing the average US citizen is GPT-4. In general, all models, except for Llama-2, align with the moral schema of a young Western liberal with a high level of education, engaged in social causes, and with a great openness to experience, empathy, and compassion.
For all models, we identified a distinctive moral profile characterized by high scores in Harm/Care and Fairness/Reciprocity, in contrast with lower values in Ingroup/Loyalty, Authority/Respect, and Purity/Sanctity. This configuration suggests a moral orientation towards empathy, compassion, and equity. Other authors have reported a notable correlation between this moral stance and a liberal or progressive political orientation (24), where individuals with this moral profile showed a preference for ideologies that emphasize social justice, welfare, and equality rights. Additionally, from a demographic perspective, this profile was predominantly found among younger individuals, particularly those in late adolescence to early adulthood, and was more prevalent among those with higher levels of education (37). Specifically, participants from Western, Educated, Industrialized, Rich, and Democratic (WEIRD) societies were more likely to exhibit this moral configuration (38). Regarding personality, it has been seen (39) that high levels of openness to experience, empathy, and compassion were significantly associated with these moral preferences. These traits highlight an individual’s inclination towards concern for the well-being of others and a commitment to justice and equity. Consistent with these moral values, individuals fitting this profile were also more actively engaged in social causes (40), such as environmentalism, human rights advocacy, and animal welfare, reflecting their moral priorities through concrete actions.
Overall, these results suggest that all models align with the moral schema of a young Western liberal with a high level of education, engaged in social causes, and with a great openness to experience, empathy, and compassion. The model that best fits this outlined profile is Claude-3-Sonnet; on the other hand, GPT-4 is most aligned with an average American citizen; the model closest to an average stance among American conservatives is Llama-2, although with significant variability. However, there is significant variability among the models (Methods - Statistical tests): there is a stark difference between open and proprietary models in the first two foundations, which is in line with results from the ethical dilemmas section.
0.3 SARA: Similarity-based Activation Steering with Repulsion and Attraction
Mechanistic interpretability (MI) is an emerging field within AI research that aims to demystify the internal workings of neural networks (41, 42), particularly large language models (LLMs). This domain focuses on understanding how these models process information and make decisions by dissecting their internal components, such as neurons, layers, and activation patterns. By revealing the underlying mechanisms, researchers can gain insights into why models behave the way they do, which is crucial for enhancing their transparency, reliability, and alignment with human values.
One of the core objectives of MI is to decode the high-dimensional representations learned by models during training. A recently developed and promising approach is that of activation patching, which takes inspiration from neuroscience (patch-clamp). Instead of directly zeroing activations (this is known as ablation (43)), patching consists in replacing model activations in a targeted way; see (44) for understanding how to best do this. A particular way in which activation patching can be applied is by shifting model activations into a particular direction of interest; this technique is known as activation steering (45, 46). The primary goal is to modify the model’s behavior in a controlled manner without altering its underlying architecture or training data. Activation steering builds on the principles of MI by leveraging the insights gained from dissecting the model’s internals (43, 47). By understanding how specific neurons and layers contribute to particular behaviors, researchers can develop techniques to modulate these components to achieve desired outcomes. This approach can be seen as a causal intervention that operates at the level of individual activations, offering a more granular and precise method of control compared to other intervention methods (such as attribution patching (48)).
In our approach, we build on the work of (46) and (49), who propose steering the activation of hidden units in response to a prompt (noted by $\mathbf{v}_{prompt}$) by directly adding activations from another template prompt $\mathbf{v}_{target}$. Additionally, they suggest subtracting activations from a third prompt $\mathbf{v}_{away}$, thereby shifting the updated activations towards the first template vector and away from the second. As is already suggested in (46), this general class of techniques work above the token level and are more general than prompt engineering. However, a potential limitation of the previous techniques is the assumption of feature linearity and independence; while it has been recently shown (50) that some small models such as Othello-GPT do show that linearity can be safely assumed, it remains to be shown whether this is a general property for other (particularly larger) models. Moreover, techniques like Activation Addition (46) are not activation-specific; they just shift all activations homogeneously, regardless of how similar they were to the desired target vector (or to the one to be repelled), to begin with.
To address these issues, we propose to adjust the model’s behavior by enhancing or suppressing specific activation patterns. Our method, Similarity-based Activation steering with Repulsion and Attraction (SARA), fine-tunes the activations of neurons in response to a given prompt (with a corresponding activation matrix $\mathbf{A}_{prompt}$) to be more similar to those activations in another prompt ($\mathbf{A}_{align}$) while being less similar to those in a third prompt ($\mathbf{A}_{repel}$). For further details on how this method is implemented, see Methods - Activation Steering: SARA.
In order to test how effective SARA was, we used Gemma-2B and compared how its unsteered and steered responses differed when addressing a particular dilemma of the previous section (criminal parent dilemma). The prompts we used to steer the model were: $x_{duties}=$ Only moral duties matter to make a moral decision, regardless of the consequences. and $x_{conseq}=$ Only consequences matter to make a moral decision, regardless of the moral duties. We steered the model response in two different conceptual directions: Kantian-steering ($x_{duties}$ as the target vector, $x_{conseq}$ as the repelled one) and Utilitarian-steering (flipping the roles of $x_{duties}$ and $x_{conseq}$). Example results can be seen in Fig. 3B, when we intervene on activations within layer $14$. Crucially, the model choice for the dilemma was unchanged (i.e., "the parent should be reported"), but the reasoning was indeed changed (valuing consequences or duties more or less, respectively). We systematically intervened on each layer (Gemma-2B has $18$ layers in total), to check how the intervention works at different processing stages and sampled 5 times, keeping fixed the temperature at 0.8080.80.8, as in Ollama, for easier comparison with previous sections. As we can see from the pooled results (Fig. 3C), SARA is effective at steering model responses in different conceptual directions (i.e., utilitarist steerings make the model respond with more utilitiarian-aligned reasonings and similarly for Kantian-steerings). If we split these results over different layers (early: layers 0-5; mid: layers 6-11; late: layers 12-17), we see (Fig. 3D) that SARA is most effective when intervening at early or late stages, whereas mid layers yield more mixed results.
Figure 3: A) Schematic of how SARA works. For details, see Activation Steering: SARA B) Example responses: unsteered (gray), Utilitarist-steered (orange) and Kantian-steered (blue). The decision (reporting the parent) is the same, the reasoning is changed. C) How different responses are classified (using the same method as in the previous section) when doing each steering. As it can be seen, SARA is effective at steering model responses in different conceptual directions. D) SARA is more effective when intervening at early or late layers, rather than at the intermediate ones.
Discussion
In this work, we have presented evidence that closed models are more aligned with an utilitarian perspective, whereas open models respond more in line with value-based ethical systems (Fig. 1B). We believe this might, in part, reflect the way in which these proprietary models are trained (i.e., demographics, value systems, etc) and the moral biases that come associated with them. Nevertheless, we also report a high variability in model response (Fig. 1D), meaning that models do not typically reliably reason following a fixed ethical perspective; this can be interpreted as low consistency or high flexibility, depending on the use case. We emphasize that it is paramount that users are aware of these biases and limitations, given that these systems are already permeating society in multiple domains: from using them as sources of entertainment to enhancing their professional output, through companies offering AI-supported services and products. Each of these examples comes with associated risks and pitfalls, that should be taken into account when making use of such systems (particularly if no humans are in the loop).
One major challenge with implementing utilitarian systems is their need to precisely predict the consequences that inform their decisions. These consequences are inherently influenced by the systems’ own actions, creating a complex feedback loop that is difficult to manage. Moreover, we argue that this issue mirrors broader concerns in artificial intelligence, particularly regarding superintelligent systems. Recent studies (51) highlight the intrinsic limitations in developing computational superintelligence. These studies show that it’s theoretically impossible to design a superintelligence with a control strategy that both prevents harm from others and ensures it cannot itself become a source of harm. This dilemma is akin to the "halting problem" in computability theory, making the "harming problem" similarly undecidable (51, 52). Thus, we conjecture that utilitarian systems cannot be said to be a priori safe, given that its code of conduct would yield an undecidable course of action. Note that if, we were to relax the perfect prediction constraint, we would then need to be open for deviations between the predicted consequences and the real ones, thus opening the door to blurry courses of actions due to misspecified targets.
When making use of the Moral Foundations Questionnaire (MFQ), we report a consistend trend across models (except for Llama-2): a distinctive moral profile characterized by high scores in Harm/Care and Fairness/Reciprocity, in contrast with lower values in Ingroup/Loyalty, Authority/Respect, and Purity/Sanctity (Fig. 2). This moral profile when answering the MFQ suggests that all models align with the moral schema of a young Western liberal with a high level of education, engaged in social causes, and with a great openness to experience, empathy, and compassion. Note that this demographic profiling is consistent with the previous hypothesis: that the moral biases and preferences of those likely designing/training these LLMs is partially leaked to the consumer-ready models.
For the last part of this work, we put forward a new method for causal intervention in LLMs: Similarity-based Activation steering with Repulsion and Attraction (SARA). We believe that SARA’s main added value comes from different key points: 1) it is designed to operate at the prompt level, therefore lowering the technical threshold needed to implement it; 2) it operates in the high-dimensional activation space, retaining much more richness than summary metrics; 3) it can also be thought of as an automated moderator, given that there is no human supervision involved in the process; 4) there is no need for prompt engineering to safeguard model responses; 5) there is no requirement on prompt lengths (for steering towards to and away from) having to be the same for this method to work. Nevertheless, we predict better steering performance when using reasonably-similarly-sized prompts; in our case, but there was a difference in prompt length of an order of magnitude.
We suggest that the role of activation steering and similar intervention techniques, apart from understanding how models process information, can be potentially used to fine-tune or safeguard foundational models without retraining. Specifically, we envision this as an extra safety layer that could be added right before the deployment stage, to further ensure that the model complies with expected behavior. This would be of particular interest for actors with a reduced access to computing power or technical resources that want to deploy pre-trained LLMs. Also, the lack of re-training or fine-tuning implies a lesser need of computational (and, thus, energetic) resources to achieve the safeguarding.
Finally, we believe it is crucial that the AI Safety field starts pivoting towards a paradigm in which there are richer performance characterizations - rather than optimizing models for certain benchmarks, which also has associated risks in itself (53). In this study, we offer hints on how one might transition into such a paradigm, benefiting from the rich existing literature in other fields and embracing a mixture of quantitative and qualitative analyses.
Conclusions
We have found that, out of the models we studied, all the proprietary ones are mostly utilitarian and all of the open-weights ones align mostly with values-based ethics. Furthermore, all models we probed - except for Llama 2- have a strong liberal bias. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked.
Acknowledgements.
We want to thank Laura Bernáez Timón, Carlos Wert Carvajal and Simón Rodríguez Santana for suggestions and feedback on earlier versions of this manuscript. A.T. acknowledges support from the Margarita Salas Fellowship (Spanish Ministry of Economy) and from the the Add-On Fellowship for Interdisciplinary Life Sciences (Joachim Herz Stiftung).